OpenStudy (anonymous):

I need help with setting up/starting these problems. 1) y=ax^2+bx+c; vertex (1,4), line passes through (-1,-8). find a b & c. answers:a=-3, b=6, c=1 2) find the point on y=x-1 that is close to (4,1) answer: (2,3)

OpenStudy (anonymous):

x coordinate for vertex : -b/2a y coordinate (you plug it into the y) so -b/2a = 1 -b = 2a now your parabola will look like : y = ax^2 -2ax + c now if you put x=1 y should be 4 4 = a -2a + c c = 4 + a so your parabola : y = ax^2 -2ax + 4 + a since the parabola passes through -1 , -8 -8 = a + 2a +4 + a 4a = -12 a = -3 b= -2a = 6 c = 4+a = 1 2) distance from the line to the point : ((x-4)^2 + (x-2)^2)^0.5 derivative will be : (4x-12)/((2x^2 -12x +2x)^0.5) equate derivative to zero : 4x-12 = 0 x = 3 y = x-1 y(3) = 2 (3,2)